An electric dipole of charges 2×10-6c,-2×10-6c are separated by distance 1cm calculate the electric field due to dipole at a point on its (i) axial line 1m from its centre (ii) equatorial line 1m from its centre

Grade:12

1 Answers

Khimraj
6 years ago
dipole moment of dipole p = q*l = 2*10-6*1*10-2Cm = 2*10-8Cm
  1. electirc field at axial line is given by E = 2Kp/r3 = 2*9*109*2*10-8/13 = 360 N/C
  2. electric field at equitorial line is given by E = Kp/r3 = 180N/C.
Hope it clears. If you like answer then please approve it.